# LatentSync: Audio Conditioned Latent Diffusion Models for Lip Sync

### v1

@@ 3.3 Empirical Studies on SyncNet Convergence
Figure 5: SyncNet training curves of different batch sizes. VoxCeleb2 results, the more transparent curves represent the training set loss, while the darker curves represent the validation set loss; the same applies to the following figures. (VoxCeleb2, Dim 2048, SD residual blocks, 16 frames.)
Architecture. We attempted to replace the SyncNet’s visual and audio encoders with the U-Net encoder from SD 1.5 [41], retaining the structure of residual blocks [21] and self-attention blocks [51] in the U-Net encoder blocks. We only adjusted the downsampling factors based on the size of the input visual images and mel-spectrograms, and we removed the cross-attention blocks, as SyncNet does not require additional conditions. As shown in Fig. 6, SyncNet using the SD U-Net encoder architecture maintained both training loss and validation loss lower than those of Wav2Lip’s SyncNet [36] throughout the training process, though the difference was not substantial.
@@
Self attention. We attempted to remove the self-attention [51] blocks in the U-Net encoder, keeping only the residual blocks [21]. As shown in Fig. 7, removing the self-attention blocks results in faster convergence during the initial training phase, while achieving lower validation loss towards the end of the training. Additionally, this approach helps to reduce the model size and accelerates both inference and training speeds. We hypothesize that the reason may be that attention layers do not possess the inherent inductive biases of CNNs, such as translation equivariance and locality. Therefore, they do not generalize well without sufficient data. Similar findings and explanations can be found in [11].
Figure 7: SyncNet training curves with and without self-attention layers. (VoxCeleb2, Dim 2048, Batch size 512, 16 frames.)
@@
Figure 8: SyncNet training curves of different numbers of input frames. (VoxCeleb2, Batch size 512, SD residual blocks, Dim 2048.)
Data preprocessing. In-the-wild videos naturally contain audio-visual offsets, so it is necessary to adjust this offset to zero before inputting them into the SyncNet network. We used the official open-source version of SyncNet [9] to adjust the offset and remove videos with ${\text{Sync}}_{\text{conf}}$ below 3. Specifically, we tested adjusting the offset before and after applying affine transformation. As shown in Fig. 9, without offset adjustment, the model’s convergence is significantly impaired. Performing affine transformation before offset adjustment yields better results. We hypothesize that this may be due to affine transformation reducing data with side profiles or unusual angles, allowing the official SyncNet [9] to predict the offset more accurately.
Figure 9: SyncNet training curves of different data preprocessing methods. (VoxCeleb2, Batch size 512, SD residual blocks, Dim 2048, 16 frames.)
Discussions. Based on the ablation studies above, we found that batch size, number of input frames, and data preprocessing method significantly impact SyncNet convergence. The model architecture can also affect convergence, though less noticeably. We identified the optimal settings for a SyncNet with the 256 $\times$ 256 input size: batch size of 1024, 16 frames, SD residual blocks for both visual and audio encoders, embedding dimension of 2048, and adjusting the audio-visual offset after affine transformation. We train a SyncNet on VoxCeleb2 [10] and test it on HDTF [62], which is an out-of-distribution experimental setup. The validation loss on VoxCeleb2 reaches around 0.18 and the accuracy on HDTF achieves 94%, which significantly surpasses the previous SOTA result 91% [36, 20]. For more ablation studies, please refer to Appendix C.
@@ 4 Experiments
FID $\downarrow$ SSIM $\uparrow$ ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$
Wav2Lip [36] 12.5 0.70 8.2 304.35 10.8 0.71 7.0 257.85
VideoReTalking [8] 9.5 0.75 7.5 270.56 7.5 0.77 6.4 215.67
DINet [63] 8.02 0.79 7.7 208.93 6.3 0.80 6.7 110.75
MuseTalk [61] 9.35 0.74 6.8 246.75 7.1 0.80 5.9 203.43
LatentSync (Ours) 7.03 0.79 8.9 192.74 5.6 0.81 7.3 124.38
@@ 4.1 Experimental Settings
Evaluation metrics. We evaluate our method in three aspects: (1) Visual quality. We use SSIM [53] in the reconstruction setting and FID [22] in the cross generation setting to assess visual quality. Following [34], the reconstruction setting refers to using the same audio as the input video, while the cross generation setting refers to using audio different from the input video. (2) Lip-sync accuracy. We use the confidence score of SyncNet (${\text{Sync}}_{\text{conf}}$) [9], which we found to align closely with visual assessments. (3) Temporal consistency. We adopt the widely used FVD metric [50].
@@ 4.2 Comparisons
Comparison methods. We selected several SOTA methods that provide open-source inference code and checkpoints for comparison. Wav2Lip [36] is the classic lip-sync method, introducing the idea of using a pretrained SyncNet for supervision instead of a lip-sync discriminator [27]. VideoReTalking [8] divides the lip-sync process into three steps to improve the results. DINet [63] proposed performing deformations on feature maps to achieve high-quality lip sync. MuseTalk [61] utilizes the architecture of Stable Diffusion for inpainting, but it is not based on the diffusion model framework and appears more like a GAN-based approach.
Quantitative comparisons. As shown in Tab. 1, the lip-sync accuracy of our method significantly surpasses that of other methods. We believe this may be attributed to our 94% accuracy SyncNet, as well as the audio cross-attention layers, which may better capture the relationship between audio and lip movements. In terms of visual quality, our method outperforms others, likely due to the capabilities of the Stable Diffusion model. Owing to the incorporation of TREPA, our FVD score is also superior to other methods.
@@
LatentSync w/o SyncNet 1.6 260.37
LatentSync + latent space SyncNet 7.9 240.45
LatentSync + pixel space SyncNet 8.9 192.74
@@
Qualitative comparisons. We run two cases in the cross generation setting. According to Fig. 10, Wav2Lip has excellent lip-sync accuracy, but the generated videos are very blurry. VideoReTalking shows strange artifacts. DINet’s generated box is obvious, and its mouth movements are stiff. MuseTalk does not preserve facial features well, the man’s beard becomes sparse. In contrast, our method excels in both clarity and identity preservation, even the mole on the woman’s face was preserved. Furthermore, it hardly shows generated box due to the smooth shape of fixed mask.
Figure 10: Qualitative comparisons with SOTA lip-sync methods. We run two cases in the cross generation setting [34]. The first row demonstrates the original input video, and the second row is the video from which we extracted the audio as input, the video can be regarded as the target lip movements. Rows 3 $\sim$ 7 display the lip-synced videos. (All the photorealistic portrait images in this paper are from licensed models.)
Figure 11: SyncNet training curves of different visual input spaces. Here we use the open-sourced pretrained VAE from Stability AI [1] for encoding. (VoxCeleb2, Batch size 512, SD residual blocks, Dim 1024, 16 frames.)
@@ 4.3 Ablation studies
The supervision effect of SyncNet. As shown in Tab. 2, when the SyncNet supervision is not added, the trained diffusion models have no lipsync ability at all. In fact, we found that other end-to-end lipsync methods [36, 34] exhibit similar phenomena. As for supervision in two different spaces, diffusion models under pixel space SyncNet supervision perform better in terms of lip-sync accuracy and temporal consistency. This is likely due to the poor convergence of latent space SyncNet supervision (as shown in Fig. 11), which is quite reasonable since the input to the latent space SyncNet is the compressed latents obtained from VAE encoding, and some lip information may already be lost.
In addition, we found that improvements in lip-sync accuracy were accompanied by an increase in temporal consistency. This is because the temporal consistency of our model is provided by the audio window; as the model becomes better at utilizing audio information, temporal consistency also improves accordingly.
Different methods for improving temporal consistency. According to Tab. 3, the temporal layer significantly improves temporal consistency but also noticeably harms lip-sync accuracy. Our TREPA achieves balanced improvements in both lip-sync accuracy and temporal consistency without evident drawbacks. Considering the importance of lip-sync accuracy, TREPA is obviously the better approach.
Method ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$
LatentSync 8.2 234.35
LatentSync + Temporal layer [19] 4.7 174.56
LatentSync + TREPA 8.9 192.74
Table 3: The ablation studies of different methods for improving temporal consistency. HDTF results.
In addition, TREPA can also improve Wav2Lip according to Tab. 4, which suggests that our method has general applicability in lip-sync tasks.
Method ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$
Wav2Lip 8.2 304.35
Wav2Lip + TREPA 8.6 264.56
Table 4: The ablation studies of Wav2Lip with TREPA. HDTF results.
@@ 5 Conclusion
We introduced a new lip sync framework named LatentSync that addresses the problems of traditional diffusion-based lip sync methods: (1) The low sampling speed and inability to generate high-resolution videos of pixel space diffusion. (2) The information loss in two-stage generation methods. In addition, we proposed TREPA to improve temporal consistency while preserving the lip-sync accuracy, which mitigates the problem of video flickering in diffusion-based lip sync methods. Furthermore, we analyzed the notorious SyncNet convergence issue and conducted comprehensive ablation studies to identify several key influencing factors.
@@
Appendix A The Differences between Lip Sync and Audio-driven Portrait Animation
Many people may confuse lip sync with audio-driven portrait animation. These two tasks have some similarities but are actually completely different tasks. Their main difference lies in the input:
  • The inputs of Lip Sync: Video + Audio
  • The inputs of Audio-driven Portrait Animation: Image + Audio
This leads to differences in their overall frameworks. Lip sync is more like a video-to-video editing framework, which needs to keep areas other than the mouth the same as in the input video. Audio-driven portrait animation is more like an image-to-video animation framework, which can change the head’s movement and even facial expressions. The differences in the overall frameworks result in substantial differences in their training and inference methods. We list some methods for the two different tasks to help readers better distinguish between them:
  • Lip Sync: Wav2Lip[36], Diff2Lip [34], StyleSync [18], MuseTalk [61], MyTalk [57]
  • Audio-driven Portrait Animation: EMO [49], Hallo [54], EchoMimic [7], VASA-1 [55], DreamTalk [31], SadTalker [60]
Although there are already some audio-driven portrait animation methods based on Stable Diffusion [41], such as EMO [49] and Hallo [54], it is hard to adapt them for the lip sync task. For example, the temporal layer is very effective in these methods, but we found that it is not suitable for the lip sync task. To the best of our knowledge, the methods closest to ours are Diff2Lip [34] and MyTalk [57], which are pixel space diffusion and two-stage generation methods.
Appendix B SyncNet Training Loss Proof
According to the classic framework of SyncNet [36, 9], we randomly provide SyncNet with positive and negative samples with a 50% probability during the training stage. The output of SyncNet is a probability distribution of whether the sample is positive or negative. We define $p(x=1)$ as the probability that the sample is positive, and $p(x=0)$ as the probability that the sample is negative. Let $p(x)$ be the true probability distribution and $q(x)$ the predicted probability distribution. We plot some scatter charts to observe the changes in the prediction probabilities of a non-converging SyncNet during the training process.
@@ Appendix B SyncNet Training Loss Proof
Appendix C Additional SyncNet Ablation Studies
C.1 Embedding dimension
The dimensions of the audio and visual embeddings in SyncNet should be carefully calibrated to a moderate size. As illustrated in Fig. 13, embeddings with smaller dimensions (e.g., 512) result in representations that fail to capture sufficient semantic information, while larger dimensions (e.g., 4096 or 6144) lead to sparse representations, thereby impeding model convergence. Through empirical analysis, we identified that an optimal embedding dimension of 2048 is suitable for images with an input resolution of 256 $\times$ 256 pixels. For images of varying resolutions, the embedding dimensions should be appropriately scaled to ensure optimal performance.
Figure 13: SyncNet training curves of different embedding dimensions. (VoxCeleb2, Batch size 512, SD residual blocks, 16 frames.)

### v2

@@ 4 Empirical Studies on SyncNet Convergence
Figure 6: SyncNet training curves of different batch sizes. VoxCeleb2 results, the more transparent curves represent the training set loss, while the darker curves represent the validation set loss; the same applies to the following figures. (VoxCeleb2, Dim 2048, StbleSyncNet arch, 16 frames.)
Architecture. We redesign the SyncNet’s visual and audio encoders with the U-Net encoder from Stable Diffusion 1.5 [33], retaining the structure of residual blocks [17] and self-attention blocks [43] in the U-Net encoder blocks. We only adjusted the downsampling factors based on the size of the input visual images and mel-spectrograms, and we removed the cross-attention blocks, as SyncNet does not require additional conditions. We refer to the SyncNet with this modified architecture as StableSyncNet. As shown in Fig. 7, StableSyncNet maintained both training loss and validation loss lower than those of Wav2Lip’s SyncNet [29] throughout the training process.
@@
Embedding dimension. As illustrated in Fig. 8, embeddings with smaller dimensions (e.g., 512) result in representations that fail to capture sufficient semantic information, while larger dimensions (e.g., 4096 or 6144) lead to sparse representations, thereby impeding model convergence. Se identified that an optimal embedding dimension of 2048 is suitable for images with an input resolution of 256 $\times$ 256 pixels.
Figure 8: SyncNet training curves of different embedding dimensions. (VoxCeleb2, Batch size 512, StableSyncNet arch, 16 frames.)
@@
Figure 9: SyncNet training curves of different numbers of input frames. (VoxCeleb2, Batch size 512, StableSyncNet arch, Dim 2048.)
Data preprocessing. In-the-wild videos naturally contain audio-visual offsets, it is necessary to adjust this offset to zero before inputting them into the SyncNet network. We used the official open-source version of pretrained SyncNet [9] to adjust the offset and remove videos with ${\text{Sync}}_{\text{conf}}$ below 3. Specifically, we evaluated adjusting the offset before and after applying affine transformation. As shown in Fig. 10, without offset adjustment, the model’s convergence is significantly impaired. Performing affine transformation before offset adjustment yields better results. This may be due to that affine transformation reducing data with side profiles or unusual angles, allowing the pretrained SyncNet [9] to predict the offset more accurately.
Figure 10: SyncNet training curves of different data preprocessing methods. (VoxCeleb2, Batch size 512, StableSyncNet arch, Dim 2048, 16 frames.)
Discussions. Based on the ablation studies above, we found that batch size, number of input frames, and data preprocessing method are the primary factors affecting SyncNet convergence. We identified the optimal settings for a StableSyncNet with the 256 $\times$ 256 input size: batch size of 1024, 16 frames, SD U-Net encoder adapted for both visual and audio encoders, embedding dimension of 2048, and adjusting the audio-visual offset after affine transformation. We train a StableSyncNet on VoxCeleb2 [10] and test it on HDTF [51], which is an out-of-distribution experimental setup. The validation loss on VoxCeleb2 reaches around 0.18 and the accuracy on HDTF achieves 94%, which significantly surpasses the previous SOTA result 91% [29, 16].
@@ 5 Experiments
FID $\downarrow$ SSIM $\uparrow$ ${\text{Sync}}_{\text{conf}}$uparrow LMD $\downarrow$ FVD $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ ${\text{Sync}}_{\text{conf}}$uparrow LMD $\downarrow$ FVD $\downarrow$
Wav2Lip [29] 12.5 0.70 8.2 0.34 304.35 10.8 0.71 7.0 0.53 257.85
VideoReTalking [8] 9.5 0.75 7.5 0.49 270.56 7.5 0.77 6.4 0.60 215.67
Diff2Lip [28] 10.3 0.72 7.9 0.36 260.45 9.8 0.73 6.9 0.54 210.45
MuseTalk [50] 9.35 0.74 6.8 0.56 246.75 7.1 0.80 5.9 0.64 203.43
LatentSync (Ours) 7.22 0.79 8.9 0.30 162.74 5.7 0.81 7.3 0.51 123.27
@@
Figure 11: Qualitative comparisons with SOTA lip-sync methods. We run two cases in the cross generation setting [28]. The first row demonstrates the original input video, and the second row is the video from which we extracted the audio as input, the video can be regarded as the target lip movements. Rows 3 $\sim$ 7 display the lip-synced videos. (All the photorealistic portrait images in this paper are from contracted models.)
@@ 5.1 Experimental Settings
Evaluation metrics. We evaluate our method in three aspects: (1) Visual quality. We use SSIM [45] in the reconstruction setting and FID [18] in the cross generation setting to assess visual quality. Following [28], the reconstruction setting refers to using the same audio as the input video, while the cross generation setting refers to using audio different from the input video. (2) Lip-sync accuracy. We use the confidence score of SyncNet (${\text{Sync}}_{\text{conf}}$) [9] and the landmark distances around the mouth (LMD) [14]. We found that ${\text{Sync}}_{\text{conf}}$ aligns closely with visual assessments. (3) Temporal consistency. We adopt the widely used FVD metric [42].
@@ 5.2 Comparisons
Comparison methods. We selected several SOTA methods that provide open-source inference code and checkpoints for comparison. Wav2Lip [29] is the classic lip-sync method, introducing the idea of using a pretrained SyncNet for supervision instead of a lip-sync discriminator [24]. VideoReTalking [8] divides the lip-sync process into three steps to improve the results. Diff2Lip [28] utilized pixel-space diffusion models to achieve generalized lip sync. MuseTalk [50] utilizes the architecture of Stable Diffusion for inpainting, but it is not based on the diffusion model framework and appears more like a GAN-based approach.
Quantitative comparisons. As shown in Tab. 1, the lip-sync accuracy of our method significantly surpasses that of other methods. This is attributed to our 94% accuracy StbaleSyncNet, as well as the audio cross-attention layers in U-Net, which better captures the relationship between audio and lip movements. In terms of visual quality, our method outperforms others, likely due to the powerful capabilities of the Stable Diffusion model. Owing to temporal layer [15] and the incorporation of TREPA, our FVD score is also superior to other methods.
@@
LatentSync w/o SyncNet 4.6 220.37
LatentSync + latent space SyncNet 7.9 180.45
LatentSync + pixel space SyncNet 8.9 162.74
@@
Qualitative comparisons. We run two cases in the cross generation setting. According to Fig. 11, Wav2Lip has excellent lip-sync accuracy, but the generated videos are very blurry. VideoReTalking shows strange artifacts. Diff2Lip is limited by pixel-space diffusion models and can only generate low-resolution videos, resulting in noticeable blurriness. MuseTalk does not preserve facial features well, the man’s beard becomes sparse. In contrast, our method excels in both clarity and identity preservation, even the mole on the woman’s face was preserved. Furthermore, it hardly shows generated box due to the smooth shape of fixed mask.
Figure 12: SyncNet training curves of different visual input spaces. Here we use the open-sourced pretrained VAE from Stability AI [1] for encoding. (VoxCeleb2, Batch size 512, StableSyncNet arch, Dim 1024, 16 frames.)
@@ 5.3 Ablation studies
The effectiveness of SyncNet supervision. As shown in Tab. 2, when the SyncNet supervision is not added, The lip-sync performance of the trained diffusion models is significantly poor. In fact, we found that other end-to-end lipsync methods [29, 28] exhibit similar phenomena. As for supervision in two different spaces, diffusion models under pixel space SyncNet supervision perform better in terms of lip-sync accuracy and temporal consistency. This is likely due to the poor convergence of latent space SyncNet (as shown in Fig. 12), which is reasonable since the input to the latent space SyncNet is the compressed latents obtained from VAE encoding, and some lip information may already be lost.
In addition, we found that improvements in lip-sync accuracy were accompanied by an increase in temporal consistency. This is because the audio window inherently encapsulates rich temporal information. Enhanced utilization of audio information by the model also leads to the improved temporal consistency.
The effectiveness of TREPA. According to Tab. 3, both the temporal consistency and visual quality improve after incorporating TREPA. This improvement may be attributed to that the robust representations extracted by VideoMAE-v2 [44] inherently encapsulate both visual and temporal information.
Method FID $\downarrow$ SSIM $\uparrow$ FVD $\downarrow$
LatentSync 7.71 0.77 176.35
LatentSync + TREPA 7.22 0.79 162.74
Table 3: The ablation studies of TREPA. HDTF results.
@@ 6 Conclusion
We introduced LatentSync, the first lip-sync method based on audio-conditioned LDMs, which addresses the problems of traditional diffusion-based lip-sync methods: (1) The low sampling speed and inability to generate high-resolution videos of pixel space diffusion. (2) The information loss in two-stage generation methods. We identified the shortcut learning problem in lip-sync task and explored different approaches to incorporate SyncNet supervision to solve this problem. We conducted comprehensive ablation studies to find out several key factors influencing SyncNet converge. In addition, we proposed TREPA to further improve the temporal consistency of our method.
@@
Appendix A SyncNet Training Loss Proof
According to the classic training framework of SyncNet [29, 9], we randomly provide SyncNet with positive and negative samples with a 50% probability during the training stage. The output of SyncNet is a probability distribution of whether the sample is positive or negative. We define $p(x=1)$ as the probability that the sample is positive, and $p(x=0)$ as the probability that the sample is negative. Let $p(x)$ be the true probability distribution and $q(x)$ the predicted probability distribution. We plot some scatter charts to observe the changes in the prediction probabilities of a non-converging SyncNet during the training process.